ESPN currently pays the NFL $1.1 million per year for eight years for the right to exclusively televise Monday Night Football. What is the net present value of this investment if the parent Disney Company has an opportunity interest rate equal to its cost of capital of 9 percent. Fox and CBS agreed to pay $712 million and $622 million respectively for six years to televise Sunday afternoon NFC game. What was that worth?  
Answer #1

ANSWER:

i = 9% and n = 8 years

ESPN paid $1.1 billion

Net present value = money paid by ESPN(p/a,i,n)

NPV = 1,100,000,000(p/a,9%,8)

NPV = 1,100,000,000 * 5.535

NPV = 6,088,500,000

so the net present value of ESPN is $6,088,500,000

Fox paid = $712,000,000

CBS paid = $622,000,000

I = 9% AND N = 6 YEARS

NPV of fox= Fox paid(p/a,i,n)

NPV of fox= 712,000,000(p/a,9%,6)

NPV of fox = 712,000,000 * 4.486

NPV of fox = 3,194,032,000

NPV of cbs = CBS paid(p/a,i,n)

NPV of cbs = 622,000,000(p/a,9%,6)

NPV of cbs = 622,000,000 * 4.486

NPV of cbs = 2,790,292,000

total worth = npv of fox + npv of cbs

total worth = 3,194,032,000 + 2,790,292,000

total worth = 5,984,324,000

so the total worth is $5,984,324,000